OpenStudy (lgbasallote):

how to integrate \[\large \int \frac{(4v^2 + 8v - 1)dv}{4v^3 + 4v^2 + v + 1}\]

OpenStudy (lgbasallote):

i just need a hint :D

OpenStudy (saifoo.khan):

1+1 = 2

OpenStudy (lgbasallote):

hmm makes sense

OpenStudy (lgbasallote):

thanks

OpenStudy (saifoo.khan):

Sure it does. ;)

OpenStudy (anonymous):

Have you tried partial fraction decomposition?

OpenStudy (anonymous):

Try factoring the denominator

OpenStudy (lgbasallote):

exactly...how??

OpenStudy (lgbasallote):

i dont get your point...

OpenStudy (lgbasallote):

=_=

OpenStudy (anonymous):

sorry.. i've been on medication..@!!!!

OpenStudy (lgbasallote):

lol

OpenStudy (anonymous):

Start with \[4v^3+4v^2+v+1\]\[4v^2(v+1)+(v+1)\]\[(4v^2+1)(v+1)\] Then do a separation of variables.

OpenStudy (lgbasallote):

makes sense

OpenStudy (anonymous):

Oops, meant partial fractions.

OpenStudy (lgbasallote):

is it a = -40/3 b =8 c = 13/3

OpenStudy (anonymous):

These are the equations that I got, and I don't think they yield the same solutions. I may have done a step incorrectly though...

OpenStudy (lgbasallote):

i see where i went wrong thanks!

OpenStudy (anonymous):

Sure thing :)

OpenStudy (lgbasallote):

a = 8 b = 0 c = -1

OpenStudy (anonymous):

Yep, that's what I got, too.

OpenStudy (lgbasallote):

yay